# Graph Flow Matching: Enhancing Image Generation with Neighbor-Aware Flow Fields

### v2

@@ 2.5 Latent-Space Generative Modeling
Generating high-resolution images directly in pixel space is computationally expensive and often burdens generative models with the need to reconstruct low-level details that are perceptually insignificant. Latent-space generative modeling mitigates this challenge by performing generation in a lower-dimensional, semantically meaningful space. This approach, popularized by models like Stable Diffusion (StableDiffusion_rombach2022high, ), uses a two-stage process: a VAE first compresses images into latent representations, and a generative model then operates within this latent space to model data distributions.
Formally, given a pre-trained VAE with encoder $E:\mathcal{X}\rightarrow\mathcal{Z}$ and decoder $D:\mathcal{Z}\rightarrow\mathcal{X}$, latent-space generative modeling learns to transform a prior distribution (e.g., Gaussian noise) into the distribution of encoded data points in $\mathcal{Z}$. The decoder $D$ then reconstructs image samples from this latent trajectory. This formulation provides several critical advantages:
@@
Building on this design, the Latent Flow Matching (LFM) framework (dao2023flow, ) extends flow matching to this perceptually aligned latent space. Rather than learning stochastic score-based dynamics as in diffusion models, LFM learns a deterministic velocity field that transports noise into encoded data samples by solving an ODE in latent space. This enables faster inference, fewer function evaluations, and simplified training compared to pixel-space diffusion.
For GFM to be effective, it must connect “neighboring” points to a graph. An embedded space is, therefore, natural to us as it defines a distance in latent space, which is easy to compute and is meaningful for complex data. Therefore, we build GFM directly on this latent space setting. We adopt the same VAE architecture as LFM—namely, the Stable Diffusion VAE—which maps $256\times 256$ RGB images into $32\times 32\times 4$ latent tensors. We retain all components of the LFM pipeline, including the constant-velocity transport plan, flow matching loss, and ODE solver. Our contribution lies exclusively in augmenting the velocity field $v_{\theta}(\mathbf{z},t)$ with a neighbor-aware graph based correction term. While there are many possible architectures for this part, here, we experimented with two architectures discussed below (Section 3).
@@ 3 Graph Flow Matching
We propose Graph Flow Matching (GFM), a modular enhancement to flow matching networks that integrates local neighborhood structure through a lightweight graph module. GFM operates under a reaction–diffusion framework and is model-agnostic, where the velocity field is decomposed into a standard reaction term which could be any off-the-shelf pointwise acting flow network, and a graph-based message passing term (the so-called “diffusion” term that we propose).
@@ 3.1 Reaction–Diffusion Decomposition
$v_{\theta}(\mathbf{x},t)=v_{\text{react}}(\mathbf{x},t)+v_{\text{diff}}(\mathbf{x},t;\,\mathcal{N}(\mathbf{x},t)),$ (7)
where $v_{\text{react}}$ is any off-the-shelf flow matching architecture, and $v_{\text{diff}}$ is a graph-based correction term informed by a neighborhood $\mathcal{N}(\mathbf{x},t)$ of surrounding samples at time $t$.
Graph Generation. In our formulation, each image represents a node on a graph, we therefore compute the graph connectivity using an attention mechanism applied to the latent representations of the images.
Reaction Term: Standard Flow Networks. The reaction term $v_{\text{react}}(\mathbf{x},t)$ can be any off-the-shelf flow matching network that predicts velocity from spatial and temporal coordinates. It typically takes the form of a U-Net (UNet_ronneberger2015u, ; Dhariwal_ADM_NEURIPS2021_49ad23d1, ), Vision Transformer (VisionTransformer_dosovitskiy2020image, ; DiT_peebles2023scalable, ), or any pointwise neural interpolator. This component models pointwise transport based on global training dynamics.
In our experiments (Sec. 4), we instantiate $v_{\text{react}}$ with a UNet (ADM) Dhariwal_ADM_NEURIPS2021_49ad23d1 and a transformer model (DiT) DiT_peebles2023scalable , following the settings in (dao2023flow, ). However, GFM is compatible with any flow matching backbone and training strategy.
Diffusion Term: Neighbor-Aware Velocity Correction. While the diffusion term $v_{\text{diff}}(\mathbf{x},t;\mathcal{N}(\mathbf{x},t))$ can be implemented using any graph-based neural architecture that aggregates information from neighboring samples, here we focus on two such architectures: the Message Passing Neural Network (MPNN) MPNN_paper_gilmer2017neural and a graph transformer architecture (GPS) GPS_rampavsek2022recipe .
Figure 3: LSUN Bedroom and LSUN Church samples ($256\times 256$) generated using the same random seed by: (top) baseline DiT-L/2 dao2023flow , (middle) DiT-L/2 with MPNN-based correction, and (bottom) DiT-L/2 with GPS-based graph correction GPS_rampavsek2022recipe . GFM variants generate more complete spatial structures and sharper boundaries compared to the baseline model.
MPNN Architecture. The first diffusion term we used is a custom MPNN architecture where we use a graph gradient ${\bf G}$ (sometimes referred to as an incidence matrix) PDE_GCN_NEURIPS2021_1f9f9d8f to compute the difference between node features. This yields an edge-based quantity, or edge features. We then apply a nonlinearity to the edge features and aggregate them back to the node using the transpose of the incidence matrix. We then apply a non-linear network to the result. The network (Equation 7) can thus be summarized by the following ODE
$v_{\theta}(\mathbf{x}_{t},t)=\frac{d\mathbf{x}_{t}}{dt}=-\,N_{\theta}^{(2)}\left({\bf G}^{\top}\left[\sigma\left({\bf G}\,N_{\theta}^{(1)}(\mathbf{x}_{t},t)\right)\right],\,t\right)+R(\mathbf{x}_{t},t),$ (8)
@@
GPS Architecture. Our second instantiation of the graph correction module employs the General, Powerful, Scalable (GPS) graph transformer architecture GPS_rampavsek2022recipe . We adapt the GPS framework for flow matching as follows: (i) recasting each latent tensor in the batch as a node in a fully connected graph; (ii) integrating temporal information via learned time embeddings that are projected and concatenated with node features; (iii) incorporating random walk positional encodings (RWPE) generated at runtime using Pytorch Geometrics’s implementation; and (iv) replacing the standard GINEConv blocks with recurrent GatedGraphConv units. The GPS architecture alternates local message passing and global multi-head attention, allowing nodes to aggregate information from the entire batch. Finally, the network projects the enriched node representations back to the original latent tensor dimensions, producing the diffusion velocity correction term. The adjacency matrix of this setup is thus attention-based.
Below, we provide a lemma that theoretically reasons about the representational advantage of GFM compared with traditional pointwise flow networks. The proof can be found in the Appendix.
Lemma 3.1 (Representational Advantage of Graph Flow Matching).
Let $\mathcal{L}_{\text{FM}}(v)$ be the flow matching loss for a velocity field $v$. The Graph Flow Matching model $v_{\theta}(\mathbf{x},t)=v_{\text{react}}(\mathbf{x},t)+v_{\text{diff}}(\mathbf{x},t)$ achieves lower or equal training loss compared to the reaction-only model $v_{\text{react}}(\mathbf{x},t)$ when both are trained to minimize the same objective:
$\mathcal{L}_{\text{FM}}(v_{\theta})\leq\mathcal{L}_{\text{FM}}(v_{\text{react}})$ (9)
Furthermore, when the data exhibits local coherence, this inequality is strict.
@@ 4 Experiments
Dataset/Model Total Parameters $v_{\text{diff}}$-net FID ($\downarrow$) FID ($\downarrow$) Recall
[…]
LSUN Church
ADM dao2023flow (Baseline) 356.38 0 - 7.70 0.39
@@
LSUN Bedroom
ADM dao2023flow (Baseline) $406.40$ 00 - $7.05$ $0.39$
@@
Dataset/Model Total Parameters $v_{\text{diff}}$-net FID ($\downarrow$) FID ($\downarrow$) Recall
@@
AFHQ-Cat
ADM (Baseline) 153.10 0 7.40 7.34 0.32
ADM+MPNN $163.17$ $10.07$ $7.01$ $6.97$ $0.35$
ADM+GPS $171.23$ $18.13$ $2.92$ $4.63$ $0.54$
DiT (Baseline) $456.80$ 00 $8.24$ $9.05$ $0.42$
DiT+MPNN $480.23$ $23.43$ $7.19$ $8.07$ $0.51$
DiT+GPS $481.25$ $24.45$ $8.21$ $8.98$ $0.42$
CelebA-HQ
ADM (Baseline) $153.10$ 00 $7.54$ $6.80$ $0.50$
ADM+MPNN $176.53$ $23.43$ $5.33$ $6.28$ $0.56$
ADM+GPS $177.55$ $24.45$ $5.89$ $6.71$ $0.52$
@@
DiT+MPNN 480.49 23.43 5.85 6.16 0.57
DiT+GPS $481.25$ $24.45$ $6.09$ $6.18$ $0.59$
We evaluate Graph Flow Matching (GFM) on five standard unconditional image generation benchmarks: LSUN Church, LSUN Bedroom, FFHQ, AFHQ-Cat, and CelebA-HQ, all at $256\times 256$ resolution. The goal is to assess the effect of incorporating neighbor-aware graph based correction into flow matching pipelines under consistent, state-of-the-art settings. We have added an additional ablation study in the Appendix.
Experimental Settings. To isolate the effect of the graph correction term, we retain all architecture and training settings from the latent flow matching (LFM) models of (dao2023flow, ). We therefore perform our experiments in the latent space of the pretrained VAE used by the aforementioned authors. Specifically, (i) Backbones: The $v_{\text{r}eact}(\mathbf{x}_{t},t)$ terms we use are the ADM U-Net (Dhariwal_ADM_NEURIPS2021_49ad23d1, ) and DiT Transformer (DiT_peebles2023scalable, ) (specifically, the ADM variants and DiT-L/2 variant from dao2023flow ), representing leading convolutional and attention-based flow architectures; (ii) Latent space: The Stable Diffusion VAE (StableDiffusion_rombach2022high, ), which maps $256\times 256$ RGB images into $32\times 32\times 4$ latent tensors. All models operate exclusively in this latent space. [^1] The specific pretrained VAE used by dao2023flow : https://huggingface.co/stabilityai/sd-vae-ft-mse; (iii) Training strategy: Constant-velocity flow matching loss (Equation 5), the Dormand–Prince (dopri5) ODE integrator with $rtol=atol=10^{-5}$, and unmodified training hyperparameters from LFM (dao2023flow, ).
This ensures that any performance gain arises from our graph-based diffusion term (i.e., the graph correction module $v_{\text{diff}}$) rather than hyperparameter tuning or architectural shifts. We evaluate the two implementations of $v_{\text{diff}}$ discussed in Section 3. The MPNN implementation uses neural components $N_{\theta}^{(1)}$ and $N_{\theta}^{(2)}$, for which we employ identical U-Net UNet_ronneberger2015u architectures derived from the U-Net design of PnPUNet_huang2021variational . The GPS implementation, like MPNN, implements attention-based adjacency. Full architectural and hyperparameter details are provided in the Appendix.
@@
Tables 1 and 2 report Fréchet Inception Distance (FID), recall, and parameter counts. Across all datasets and backbones, GFM consistently improves sample quality, achieving FID reductions of up to and sometimes exceeding $40\%$ relative to the base models. These improvements are achieved with only modest computational overhead—typically less than a $10\%$ increase in total parameters.
Qualitatively, GFM-enhanced models generate better-structured furniture and spatial layouts on LSUN Bedroom, sharper and more anatomically plausible faces on FFHQ, and more coherent architectural forms on LSUN Church as shown in Figures 23. The effectiveness of GFM across both convolutional (ADM) and transformer-based (DiT) models underscores its architectural generality.
To further assess the role of graph structure, we conduct an ablation in which the adjacency matrix of the GPS module is replaced with the identity matrix, thereby eliminating inter-node communication while preserving architectural and parameter parity. Note that for the MPNN module, setting the adjacency matrix to identity would lead to a zero gradient matrix ${\bf G}$. This would effectively cancel out the diffusion term, leaving only the reaction term $v_{\text{react}}$, making the MPNN architecture unsuitable for this particular ablation study. This ablation reduces GPS to a non-graph network and thus isolates the effect of the graph from the addition of a second network. Results in Table 3 show a clear degradation in FID and recall when graph connectivity is removed, confirming that performance gains stem from graph structure rather than parameter count or the addition of a second network alone. Notably, GFM provides consistent performance gains across both convolutional (ADM) and transformer-based (DiT) architectures. This underscores the generality of GFM as an architectural enhancement that is agnostic to the flow matching backbone $v_{\text{react}}$.
Table 3: Ablation study on graph structure in LSUN Church, FFHQ, and AFHQ-Cat datasets. We compare our graph-based diffusion term with the self-only diffusion baseline where the graph adjacency matrix is set to identity (Adj=I). FID scores and recall show that the graph structure improves performance.
Dataset/Model Total Parameters $v_{\text{diff}}$-net FID ($\downarrow$) FID ($\downarrow$) Recall
@@
LSUN Church
ADM+GPS $374.20$ $18.13$ $4.67$ $4.61$ $0.51$

### v3

@@ 2.5 Latent-Space Generative Modeling
Generating high-resolution images directly in pixel space is computationally expensive and often burdens generative models with the need to reconstruct low-level details that are perceptually insignificant. Latent-space generative modeling mitigates this challenge by performing generation in a lower-dimensional, semantically meaningful space. This approach, popularized by models like Stable Diffusion (StableDiffusion_rombach2022high), uses a two-stage process: a VAE first compresses images into latent representations, and a generative model then operates within this latent space to model data distributions.
Formally, given a pre-trained VAE with encoder $E:\mathcal{I}\rightarrow\mathcal{X}$ and decoder $D:\mathcal{X}\rightarrow\mathcal{I}$, latent-space generative modeling learns to transform a prior distribution (e.g., Gaussian noise) into the distribution of encoded data points in $\mathcal{X}$. Here $\mathcal{I}=[-1,1]^{H\times W\times C}$ denotes the pixel-space of [1,1][-1,1]-normalized $H\times W\times C$ images (as used by the Stable Diffusion VAE), and $\mathcal{X}\subseteq\mathbb{R}^{d}$ its latent representation. The decoder $D$ then reconstructs image samples from this latent trajectory in $\mathcal{X}$ to $\mathcal{I}$. This formulation provides several critical advantages:
@@
Building on this design, the Latent Flow Matching (LFM) framework (dao2023flow) extends flow matching to this perceptually aligned latent space. Rather than learning stochastic score-based dynamics as in diffusion models, LFM learns a deterministic velocity field that transports noise into encoded data samples by solving an ODE in latent space. This enables faster inference, fewer function evaluations, and simplified training compared to pixel-space diffusion.
For GFM to be effective, it must connect ”neighboring” points to a graph. An embedded space is, therefore, natural to us as it defines a distance in latent space, which is easy to compute and is meaningful for complex data. Therefore, we build GFM directly on this latent space setting. We adopt the same VAE architecture as LFM—namely, the Stable Diffusion VAE—which maps $256\times 256$ RGB images into $32\times 32\times 4$ latent tensors. We retain all components of the LFM pipeline, including the constant-velocity transport plan, flow matching loss, and ODE solver. Our contribution lies exclusively in augmenting the velocity field $\mathbf{v}_{\theta}(\mathbf{x},t)$ with a neighbor-aware graph based correction term. While there are many possible architectures for this part, here, we experimented with two architectures discussed below (Section 3).
@@ 3 Graph Flow Matching
We propose Graph Flow Matching (GFM), a modular enhancement to flow matching networks that integrates local neighborhood structure through a lightweight graph module. GFM operates under a reaction–diffusion framework and is model-agnostic, where the velocity field is decomposed into a standard reaction term which could be any off-the-shelf pointwise acting flow network, and a graph-based message passing term (the so-called ”diffusion” term that we propose).
@@ 3.1 Reaction–Diffusion Decomposition
$\mathbf{v}_{\theta}(\mathbf{x},t)=\mathbf{v}_{\text{react}}(\mathbf{x},t)+\mathbf{v}_{\text{diff}}(\mathbf{x},t;\,\mathcal{N}(\mathbf{x},t)),$ (7)
where $\mathbf{v}_{\text{react}}$ is any off-the-shelf flow matching architecture, and $\mathbf{v}_{\text{diff}}$ is a graph-based correction term informed by a neighborhood $\mathcal{N}(\mathbf{x},t)$ of surrounding samples at time $t$.
Graph Generation. At each intermediate timestep $t$, we treat the VAE latent code $\mathbf{x}_{t}\in\mathbb{R}^{4\times 32\times 32}$ of every sample in the minibatch as a graph node. We compute pairwise attention scores between these nodes. The attention weights form the adjacency matrix $\mathbf{A}$, so that $\mathcal{N}(\mathbf{x}_{t},t)=\{\mathbf{x}_{j}\mid\mathbf{A}_{ij}>0\}$ is exactly the set of latent codes connected to $\mathbf{x}_{t}$, with node index $i$ in the graph. In Appendix C we provide an ablation where the K-nearest-neighbors with learnable weights was used to construct the graph.
Reaction Term: Standard Flow Networks. The reaction term $v_{\text{react}}(\mathbf{x},t)$ can be any off-the-shelf flow matching network that predicts velocity from spatial and temporal coordinates. It typically takes the form of a U-Net (UNet_ronneberger2015u; Dhariwal_ADM_NEURIPS2021_49ad23d1), Vision Transformer (VisionTransformer_dosovitskiy2020image; DiT_peebles2023scalable), or any pointwise neural interpolator. This component models pointwise transport based on global training dynamics.
In our experiments (Sec. 4), we instantiate $v_{\text{react}}$ with a UNet (ADM) (Dhariwal_ADM_NEURIPS2021_49ad23d1) and a transformer model (DiT) (DiT_peebles2023scalable), following the settings in (dao2023flow). However, GFM is compatible with any flow matching backbone and training strategy.
Diffusion Term: Neighbor-Aware Velocity Correction. While the diffusion term $\mathbf{v}_{\text{diff}}(\mathbf{x},t;\mathcal{N}(\mathbf{x},t))$ can be implemented using any graph-based neural architecture that aggregates information from neighboring samples, here we focus on two such architectures: the Message Passing Neural Network (MPNN) (MPNN_paper_gilmer2017neural) and a graph transformer architecture (GPS) (GPS_rampavsek2022recipe).
MPNN Architecture. The first diffusion term we used is a custom MPNN architecture where we use a graph gradient $\bf G$} (sometimes referred to as an incidence matrix) (PDE_GCN_NEURIPS2021_1f9f9d8f) to compute the difference between node features. This yields an edge-based quantity, or edge features. We then apply a nonlinearity to the edge features and aggregate them back to the node using the transpose of the incidence matrix. We then apply a non-linear network to the result. The network (Equation 7) can thus be summarized by the following ODE
$\begin{split}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)&=\frac{d\mathbf{x}_{t}}{dt}\\[-1.29167pt] &=-\,N^{(2)}_{\!\theta}\bigl(G^{\top}[\sigma(G\,N^{(1)}_{\!\theta}(\mathbf{x}_{t},t))],\,t\bigr)+R(\mathbf{x}_{t},t)\,.\end{split}$ (8)
@@
GPS Architecture. Our second instantiation of the graph correction module employs the General, Powerful, Scalable (GPS) graph transformer architecture (GPS_rampavsek2022recipe). We adapt the GPS framework for flow matching as follows: (i) recasting each latent tensor in the batch as a node in a fully connected graph; (ii) integrating temporal information via learned time embeddings that are projected and concatenated with node features; (iii) incorporating random walk positional encodings (RWPE) generated at runtime using Pytorch Geometrics’s implementation; and (iv) replacing the standard GINEConv blocks with recurrent GatedGraphConv units. The GPS architecture alternates local message passing and global multi-head attention, allowing nodes to aggregate information from the entire batch. Finally, the network projects the enriched node representations back to the original latent tensor dimensions, producing the diffusion velocity correction term. The adjacency matrix of this setup is thus attention-based.
@@ 4 Experiments
Dataset Model Total $\mathbf{v}_{\mathit{diff}}$ FID ($\downarrow$) FID ($\downarrow$) Recall
@@
LSUN Church ADM (Baseline) 356.38 0 7.70 0.39
@@
LSUN Bedroom ADM (Baseline) 406.40 0 7.05 0.39
@@
Dataset Model Total $\mathbf{v}_{\mathit{diff}}$ FID ($\downarrow$) FID ($\downarrow$) Recall
@@
AFHQ‐Cat ADM (Baseline) 153.10 0 7.40 7.34 0.32
ADM+MPNN (Ours) 163.17 10.07 7.01 6.97 0.35
ADM+GPS (Ours) 171.23 18.13 2.92 4.63 0.54
DiT (Baseline) 456.80 0 8.24 9.05 0.42
DiT+MPNN (Ours) 480.23 23.43 7.19 8.07 0.51
DiT+GPS (Ours) 481.25 24.45 8.21 8.98 0.42
CelebA‐HQ ADM (Baseline) 153.10 0 7.54 6.80 0.50
ADM+MPNN (Ours) 176.53 23.43 5.33 6.28 0.56
ADM+GPS (Ours) 177.55 24.45 5.89 6.71 0.52
@@
DiT+MPNN (Ours) 480.49 23.43 5.85 6.16 0.57
DiT+GPS (Ours) 481.25 24.45 6.09 6.18 0.59
@@
We evaluate Graph Flow Matching (GFM) on five standard unconditional image generation benchmarks: LSUN Church, LSUN Bedroom, FFHQ, AFHQ-Cat, and CelebA-HQ, all at $256\times 256$ resolution. The goal is to assess the effect of incorporating neighbor-aware graph based correction into flow matching pipelines under consistent, state-of-the-art settings. We also conduct experiments with sparse K-nearest neighbor graphs (Appendix C), which demonstrate that GFM provides benefits even with sparse graphs.
Experimental Settings. To isolate the effect of the graph correction term, we retain all architecture and training settings from the latent flow matching (LFM) models of (dao2023flow). We therefore perform our experiments in the latent space of the pretrained VAE used by the aforementioned authors. Specifically, (i) Backbones: The $\mathbf{v}_{\text{r}eact}(\mathbf{x}_{t},t)$ terms we use are the ADM U-Net (Dhariwal_ADM_NEURIPS2021_49ad23d1) and DiT Transformer (DiT_peebles2023scalable) (specifically, the ADM variants and DiT-L/2 variant from (dao2023flow)), representing leading convolutional and attention-based flow architectures; (ii) Latent space: The Stable Diffusion VAE (StableDiffusion_rombach2022high), which maps $256\times 256$ RGB images into $32\times 32\times 4$ latent tensors. All models operate exclusively in this latent space [^1] The pretrained VAE used by (dao2023flow): https://huggingface.co/stabilityai/sd-vae-ft-mse; (iii) Training strategy: Constant-velocity flow matching loss (Equation 5), the Dormand–Prince (dopri5) ODE integrator with $rtol=atol=10^{-5}$, and unmodified training hyperparameters from LFM (dao2023flow).
This ensures that any performance gain arises from our graph-based diffusion term (i.e., the graph correction module $\mathbf{v}_{\text{diff}}$) rather than hyperparameter tuning or architectural shifts. We evaluate the two implementations of $\mathbf{v}_{\text{diff}}$ discussed in Section 3. The MPNN implementation uses neural components $N_{\theta}^{(1)}$ and $N_{\theta}^{(2)}$, for which we employ identical U-Net (UNet_ronneberger2015u) architectures derived from the U-Net design of (PnPUNet_huang2021variational). The GPS implementation, like MPNN, implements attention-based adjacency. Full architectural and hyperparameter details are provided in the Appendix.
@@
Table 1 and Table 2 report Fréchet Inception Distance (FID), recall, and parameter counts. Across all datasets and backbones, GFM consistently improves sample quality, achieving FID reductions of up to and sometimes exceeding $40\%$ relative to the base models. These improvements are achieved with only modest computational overhead—typically less than a $10\%$ increase in total parameters.
Qualitatively, GFM-enhanced models generate better-structured furniture and spatial layouts on LSUN Bedroom, sharper and more anatomically plausible faces on FFHQ, and more coherent architectural forms on LSUN Church as shown in Figures 2 and Figure 3 (in the Appendix). The effectiveness of GFM across both convolutional (ADM) and transformer-based (DiT) models underscores its architectural generality.
To further assess the role of graph structure, we conduct an ablation in which the adjacency matrix of the GPS module is replaced with the identity matrix, thereby eliminating inter-node communication while preserving architectural and parameter parity. Note that for the MPNN module, setting the adjacency matrix to identity would lead to a zero gradient matrix $\bf G$}. This would effectively cancel out the diffusion term, leaving only the reaction term $\mathbf{v}_{\text{react}}$, making the MPNN architecture unsuitable for this particular ablation study. This ablation reduces GPS to a non-graph network and thus isolates the effect of the graph from the addition of a second network. Results in Table 3 show a clear degradation in FID and recall when graph connectivity is removed, confirming that performance gains stem from graph structure rather than parameter count or the addition of a second network alone. This conclusion is further supported by our KNN ablation (Appendix C), where sparse graphs with controlled parameter counts still outperform wider baseline models, demonstrating that the benefits arise specifically from enabling inter-node communication, whether through dense or sparse connectivity, rather than simply from increased model capacity. Notably, GFM provides consistent performance gains across both convolutional (ADM) and transformer-based (DiT) architectures. This underscores the generality of GFM as an architectural enhancement that is agnostic to the flow matching backbone $\mathbf{v}_{\text{react}}$.
Dataset Model Total $\mathbf{v}_{\mathit{diff}}$ FID ($\downarrow$) FID ($\downarrow$) Recall
[…]
LSUN Church ADM+GPS 374.20 18.13 4.67 4.61 0.51
@@
Table 3: Ablation study isolating graph structure benefits in LSUN Church, FFHQ, and AFHQ-Cat. We compare our full graph augmented networks against an identical-parameter baseline where the graph adjacency matrix is set to identity (Adj=I), eliminating inter-node communication while preserving all network parameters. Total parameters include the $\mathbf{v}_{\mathit{diff}}$ parameters. Performance degradation when graph connectivity is removed (Adj=I) demonstrates that gains arise from graph structure, not just parameter count.